# Zero123-6D: Zero-shot Novel View Synthesis for RGB Category-level 6D Pose Estimation

### v1

@@ Abstract
Estimating the pose of objects through vision is essential to make robotic platforms interact with the environment. Yet, it presents many challenges, often related to the lack of flexibility and generalizability of state-of-the-art solutions. Diffusion models are a cutting-edge neural architecture transforming 2D and 3D computer vision, outlining remarkable performances in zero-shot novel-view synthesis. Such a use case is particularly intriguing for reconstructing 3D objects. However, localizing objects in unstructured environments is rather unexplored. To this end, this work presents Zero123-6D to demonstrate the utility of Diffusion Model-based novel-view-synthesizers in enhancing RGB 6D pose estimation at category-level by integrating them with feature extraction techniques. The outlined method exploits such a novel view synthesizer to expand a sparse set of RGB-only reference views for the zero-shot 6D pose estimation task. Experiments are quantitatively analyzed on the CO3D dataset, showcasing increased performance over baselines, a substantial reduction in data requirements, and the removal of the necessity of depth information.
@@ I Introduction
Given the data requirements for generalization to unseen objects, recent efforts have focused on simplifying object pose estimation through online optimization, reframing the task as relative pose estimation between a novel query object and a reference. However, existing methods still rely on full scans [2, 3], CAD models, [4, 5], or both [6] for each new instance, which limits their usability in practice.
@@
To address the aforementioned limitations, this document presents Zero123-6D: it exploits pre-trained diffusion models for zero-shot novel-view RGB synthesis of object images in the wild [11, 12, 13] to expand the available set of reference views. Except for the CAD reconstruction phase (which is performed once per category), none of the models employed in the framework requires any training phase. The overall pipeline just requires a monocular input query of an object instance and a sparse set of posed RGB images of a reference object from the same category. Poses are estimated through feature extraction with a foundation model and online pose optimization based on established feature correspondences.
Zero123-6D reduces previous data requirements and enhances applicability in diverse settings thanks to the pre-trained novel-view synthesizer. To the best of our knowledge, this is the first work that tries to employ a zero-shot diffusion model as a novel-view synthesizer to address the data availability in the context of category-level pose estimation. Even if only RGB images are exploited with no depth information, the pipeline shows increased performance with respect to the existing state of the art.
The main contributions of this work can be summarized as follows:
@@
  • a refinement technique to compensate for the lack of depth information leveraging the newly synthesized views.
@@ II-B Category-level Pose Estimation
Category-level pose estimation addresses this challenge by enhancing generalizability across multiple objects within a semantic category despite substantial intra-class variability. NOCS [25] presented a new benchmark for this task that showcased relevant improvements starting from [26, 27, 28] to more recent methods like [29, 30, 31]. Nevertheless, this class of methods does not scale well because it requires new training and a new dataset for each class, and data is typically limited [32]. To overcome such problems, OnePose [2] and OnePose++ [3] deal with novel objects by acquiring a full RGB scan of the item followed by a structure from the motion pipeline. The main limitation in this case is the need for a full scan and reconstruction for each new object, limiting its applicability in the case of efficient processing requirements. POPE [33] represents an alternative approach, which employs off-the-shelf methods [34, 10] to retrieve the camera pose between different views. Nonetheless, at least one camera-to-object estimate is required to make the method useful for robotic applications. To this end, Zero-Shot-Pose (ZSP) [7] aimed to enhance the generalizability of 6D pose estimation across multiple objects within a given semantic class. The method depicts a relative pose estimation setting but assumes depth availability for both query and reference objects.
ZSP uses one RGB-D query image and a set of real-world RGB-D views of the reference object to extract semantic feature correspondences across them with a foundation model. Then, it lifts them to 3D with the available depth maps and computes the alignment. Other than depth availability, the main drawback of ZSP is the viewpoint variability in the reference set: if not enough, it can hinder a proper rotation estimation. In light of this, the focus is on a generative approach using a diffusion model to synthesize novel views from a very sparse RGB-only reference set.
@@ III Methodology
Our zero-shot pose estimation approach encompasses three steps: first, the reference set composed of $N$ RGB views is expanded to obtain a set of $M>N$ RGB views. This is achieved by using a pre-trained view-conditioned diffusion model novel-view synthesizer, which generates RGB novel views of objects in the wild (Section III-A). All the generated views, together with their poses, are fed inside an implicit neural surface reconstructor NeuS [45] to obtain a 3D CAD model of the object. Second, the reference object and target object are matched through semantic correspondence by using DINO feature extractor [9]. Based on these correspondences, the most informative image is selected from the expanded reference sequence (M images) for pose estimation (III-B). At this point, the pose of the selected view represents a first coarse estimation of the 6D pose. This pose is used to render its correspondent depth from the reconstructed CAD model. Finally, the coarse pose is refined: the orientation and position transformation are computed between the reference and query objects using 3D reference depth information and the identified 2D query correspondences. Fig. 2 depicts the schematic overview of the approach.
@@ III-A Novel view synthesis
The diffusion-model based novel view synthesizer employed in the presented work is EscherNet [12]. Given one or more RGB views of an object $\mathbf{I}_{1:N}\in\mathbb{R}^{H\times W\times 3}$ with their camera poses $\mathbf{P}_{1:N}$ and a set of target views $\tilde{\mathbf{I}}_{1:M}\in\mathbb{R}^{H\times W\times 3}$ with their desired camera poses $\mathbf{P}_{1:M}$, EscherNet is able to generate multi RGB novel views $\tilde{\mathbf{I}}_{1:M}$ of the object in the specified poses. The view synthesis is formulated as a conditional generative problem:
@@ III-B Feature extraction and view extraction
Once the new RGB images are rendered using the diffusion model, it is possible to extract semantic features by relying on various foundation models. Among these DINO [9] has proven to be successfully suitable to extract semantic correspondences between images of different instances. This capability is particularly appealing for category-level tasks, where a semantic relationship between objects of the same class is enforced. As explored in ZSP [7], semantic features can be extracted from image $\mathbf{I_{q}}\in\mathbb{R}^{H\times W\times 3}$ and a reference $\mathbf{I_{r}}\in\mathbb{R}^{H\times W\times 3}$, with the operator $\mathbf{F}:\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{h^{\prime}\times w^{\prime}\times c}$, with $h^{\prime}$, $w^{\prime}$, $c$ height, width and channel size of the feature map.
These features can be leveraged to establish matches between a query using a soft nearest neighbor operator (SNN):
@@
where $\mathbf{F}_{j}[\mathbf{x}_{j}]$ denotes the feature $\mathbf{F}_{j}$ corrispondent to coordinate $\mathbf{x}_{j}$ and the function $\Gamma\left(\mathbf{F}_{i},\mathbf{F}_{j}\left[\mathbf{x}_{j}\right]\right)=:\bar{\mathbf{x}}_{i}$ selects the coordinate $\bar{\mathbf{x}}_{i}$ on the feature map $\mathbf{F}_{i}$ which is closest in feature space to $\mathbf{F}_{j}[\mathbf{x}_{j}]$.
@@
However, once many views are synthesized by EscherNet, NeRF-based reconstruction technique [45] is exploited to produce a 3D mesh using a ray-marching algorithm. From the mesh, it can be extracted the depth $\mathbf{D}_{r}$ aligned with the best RGB image $\tilde{\mathbf{I}}_{i}^{\ast}$ by simply rendering the CAD on the corresponding reference pose.
@@
We optimize for both rotation and translation in terms of summed reprojection residuals for matched feature points through
$\left(\mathbf{R}_{q},\mathbf{t}_{q}\right)=\operatorname*{argmin}_{\mathbf{R},\mathbf{t}}\sum_{k=1}^{K}\left\lVert\mathbf{K}_{q}\left(\mathbf{R}\mathbf{X}_{r}^{k}+\mathbf{t}\right)-\mathbf{x}_{q}^{k}\right\lVert_{2},$ (8)
where $\mathbf{R}_{q},\mathbf{t}_{q}$ are the resulting query rotation and translation respectively. This iterative refinement process leverages first-order stochastic gradients calculated with Adam [48] to gradually align feature points while optimizing for pose parameters.
@@ IV Experiments
Since the problem falls in the up-to-scale relative pose estimation between a query and one or more references, as in [7, 33, 50], the focus is on rotation error (computed as the Geodesic distance on the $\text{SO}(3)$ manifold of rotation matrices). Nevertheless, it is worth remembering that the method is capable of estimating the translation vector, which, however, must be opportunely rescaled for real-world applications.
The assessment includes median rotation error (in degrees) and accuracy. The latter is computed as the percentage of estimates that fall below a certain rotation threshold. Here, 15 and 30 degrees thresholds have been chosen, following ZSP [7]. Results are provided both per category and average over all the categories.
Additionally, qualitative results are presented on the Objectron dataset [51] to highlight the cross-dataset generalization ability of the proposed method to generalize to diverse RGB-only settings. All the experiments have been run on a single NVidia RTX3090 GPU. For each set of reference images, 50 novel views are generated as a trade-off between generation and 3D model reconstruction timing. The pose estimation optimizer runs 1000 iterations with an early stop on the loss computed as in Eq. 8. This step takes 2-4 seconds to run, depending on the above stopping criteria. The most time-intensive step is the 3D CAD generation, which, however, is carried out only once per category, so in the long run, it does not influence the real-time factor of the application.
@@ IV-A1 Comparison against depth-free approach
Since the presented approach is RGB-only, it is worth starting a quantitative evaluation against methods employing no-depth images, which generally provide a great geometric prior and coherently improved performance.
The concurrent work ZSP presented a depth-free version, based on the Essential Matrix calculation [7],[52]. The Essential Matrix embeds the relative pose information between two RGB views from a set of matches. It is typically used in stereo-vision, yet in this case has been effectively employed, by exploiting only the available sparse semantical correspondences.
@@ IV-A2 Comparison against RGB-D approach
The second mode, analyzed in Table II, encompasses the full method, where ZSP can show better performance only when enough views are available, by exploiting depth information for both query and reference. As it appears from Table II, with the diffusion-based novel-view synthesis, it is possible to substantially improve results over best view selection carried out in ZSP [7], thanks to having more images available for the matching phase. At the same time, this indicates the reduced sim-to-real gap between real and generated images, highlighting how diffusion models can be useful for this task. On the other hand, as shown in Table III, the Umeyama-RANSAC depth-based refinement method provided by ZSP increases the performance with respect to its correspondent best-view. This method requires an available and employable depth and boosts ZSP’s performance; however, it is eroded in few-view reference scenarios, as it is possible to notice a cross-over point with three views. Ultimately, Zero123-6D completes the overtake on ZSP when just a single reference view is available. When reducing from three views to one, Zero123-6D maintains a proper understanding of the 3D structure of the object. At the same time, the depth cues can no longer compensate for the small number of views, as in the previous cases. This can become relevant in fixed camera settings scenarios, where acquiring multiple views of the reference object is generally difficult.
@@ IV-B Qualitative results
This section highlights the qualitative results of the method in accordance with the aforementioned discussion about numerical performance. Fig.4 shows the feature map of the instances and the selected best-view. The best-view matching is possible thanks to the semantical information provided by DINO features. Applying Principal Component Analysis (PCA) on the feature map makes it possible to see how similar parts of the objects are depicted with similar colors. This is reflected in the quality of matches, which makes it possible to initialize the presented refiner.
@@
As further experiments, qualitative results on the Objectron [53] dataset are presented. These are particularly interesting because Objectron, unlike CO3D, does not have depth maps associated with RGB images. Hence, in this scenario, methods like ZSP [7], could not be effectively employed without a performance drop, while Zero123-6D’s method can be applied unaltered. Results presented in Fig.4 depict such cross-dataset generalization, representing a promising direction to employ objects generated through novel-view synthesis for pose estimation in more challenging setups (e.g. cluttered scenarios).

### v2

# {textblock}

@@ Abstract
Estimating the pose of objects through vision is essential to make robotic platforms interact with the environment. Yet, it presents many challenges, often related to the lack of flexibility and generalizability of state-of-the-art solutions. Diffusion models are a cutting-edge neural architecture transforming 2D and 3D computer vision, outlining remarkable performances in zero-shot novel-view synthesis. Such a use case is particularly intriguing for reconstructing 3D objects. However, localizing objects in unstructured environments is rather unexplored. To this end, this work presents Zero123-6D, the first work to demonstrate the utility of Diffusion Model-based novel-view-synthesizers in enhancing RGB 6D pose estimation at category-level, by integrating them with feature extraction techniques. Novel View Synthesis allows to obtain a coarse pose that is refined through an online optimization method introduced in this work to deal with intra-category geometric differences. In such a way, the outlined method shows reduction in data requirements, removal of the necessity of depth information in zero-shot category-level 6D pose estimation task, and increased performance, quantitatively demonstrated through experiments on the CO3D dataset.
@@ I Introduction
Given the data requirements for generalization to unseen objects, recent efforts have focused on simplifying object pose estimation through online optimization, reframing the task as relative pose estimation between a novel query object and a reference. However, existing methods still rely on full scans [2, 3], CAD models, [4, 5], or both [6, 7] for each new instance, which limits their usability in practice.
@@
To address the aforementioned limitations, this document presents Zero123-6D: it exploits pre-trained diffusion models for zero-shot novel-view RGB synthesis of object images in the wild [12, 13, 14] to expand the available set of reference views. Except for the CAD reconstruction phase (which is performed once per category), none of the models employed in the framework requires any training phase. The overall pipeline just requires a monocular input query of an object instance and a sparse set of posed RGB images of a reference object from the same category. Poses are estimated through feature extraction with a foundation model and online pose optimization based on established feature correspondences. Zero123-6D reduces previous data requirements and enhances applicability in diverse settings thanks to the pre-trained novel-view synthesizer. To the best of our knowledge, this is the first work that tries to employ a zero-shot diffusion model as a novel-view synthesizer to address the data availability in the context of category-level pose estimation. Even if only RGB images are exploited with no depth information, the pipeline shows increased performance with respect to the existing state of the art. The main contributions of this work can be summarized as follows:
@@
  • a novel refinement technique to compensate intra-class geometrical variability.
@@ II-B Category-level Pose Estimation
Category-level pose estimation addresses this challenge by enhancing generalizability across multiple objects within a semantic category despite substantial intra-class variability. NOCS [26] presented a new benchmark for this task that showcased relevant improvements starting from [27, 28, 29] to more recent methods like [30, 31, 32]. Nevertheless, this class of methods does not scale well because it requires new training and a new dataset for each class, and data is typically limited [33]. To overcome such problems, OnePose [2] and OnePose++ [3] deal with novel objects by acquiring a full RGB scan of the item followed by a structure from the motion pipeline. The main limitation in this case is the need for a full scan and reconstruction for each new object, limiting its applicability in the case of efficient processing requirements. POPE [34] represents an alternative approach, which employs off-the-shelf methods [35, 11] to retrieve the camera pose between different views. Nonetheless, at least one camera-to-object estimate is required to make the method useful for robotic applications. To this end, Zero-Shot-Pose (ZSP) [8] aimed to enhance the generalizability of 6D pose estimation across multiple objects within a given semantic class. The method depicts a relative pose estimation setting but assumes depth availability for both query and reference objects. ZSP uses one RGB-D query image and a set of real-world RGB-D views of the reference object to extract semantic feature correspondences across them with a foundation model. Then, it lifts them to 3D with the available depth maps and computes the alignment. Other than depth availability, the main drawback of ZSP is the viewpoint variability in the reference set: if not enough, it can hinder a proper rotation estimation. In light of this, the focus is on a generative approach using a diffusion model to synthesize novel views from a very sparse RGB-only reference set.
@@ III Methodology
Our zero-shot pose estimation approach encompasses three steps. First, the reference set is expanded using a zero-shot novel-view synthesizer; all the generated views and their poses are fed inside a 3D neural reconstructor to obtain a 3D CAD model of the object. Secondly, the reference object and target object are matched through semantic correspondence by using pre-trained feature extractors [10]; based on these correspondences, the most informative image is selected from the expanded reference sequence. At this point, the pose of the selected view represents the first coarse estimation of the 6D pose, which is used to render its correspondent depth from the reconstructed CAD model. Finally, the coarse pose is refined by leveraging 2D-3D correspondences. Fig. 2 depicts the schematic overview of the approach.
@@ III-A Novel view synthesis
The diffusion-model based novel view synthesizer employed in the presented work is EscherNet [13]. Given 1 or more (N) RGB views of an object $\mathbf{I}_{1:N}\in\mathbb{R}^{H\times W\times 3}$ with their camera poses $\mathbf{P}_{1:N}$ and a set of 1 or more (M) target views $\tilde{\mathbf{I}}_{1:M}\in\mathbb{R}^{H\times W\times 3}$ with their desired camera poses $\mathbf{P}_{1:M}$, EscherNet is able to generate multi RGB novel views $\tilde{\mathbf{I}}_{1:M}$ of the object in the specified poses. The view synthesis is formulated as a conditional generative problem:
@@ III-B Feature extraction and view extraction
Once the new RGB images are rendered using the diffusion model, it is possible to extract semantic features by relying on various foundation models. Among these DINO [10] has proven to be successfully suitable to extract semantic correspondences between images of different instances. This capability is particularly appealing for category-level tasks, where a semantic relationship between objects of the same class is enforced. As explored in ZSP [8], semantic features can be extracted from image $\mathbf{I_{q}}\in\mathbb{R}^{H\times W\times 3}$ and a reference $\mathbf{I_{r}}\in\mathbb{R}^{H\times W\times 3}$, with the operator $\mathbf{F}:\mathbb{R}^{H\times W\times 3}\rightarrow\mathbb{R}^{h^{\prime}\times w^{\prime}\times c}$, with $h^{\prime}$, $w^{\prime}$, $c$ height, width and channel size of the feature map. These features can be leveraged to establish matches between a query using a soft nearest neighbor operator (SNN):
@@
where $\mathbf{F}_{j}[\mathbf{x}_{j}]$ denotes the feature $\mathbf{F}_{j}$ correspondent to coordinate $\mathbf{x}_{j}$ and the function $\Gamma\left(\mathbf{F}_{i},\mathbf{F}_{j}\left[\mathbf{x}_{j}\right]\right)=:\bar{\mathbf{x}}_{i}$ selects the coordinate $\bar{\mathbf{x}}_{i}$ on the feature map $\mathbf{F}_{i}$ which is closest in feature space to $\mathbf{F}_{j}[\mathbf{x}_{j}]$.
@@
However, once many views are synthesized by EscherNet, NeuS [46] (a NeRF-based reconstruction technique) is exploited to produce a 3D mesh using a ray-marching algorithm. From the mesh, it can be extracted the depth $\mathbf{D}_{r}$ aligned with the best RGB image $\tilde{\mathbf{I}}_{i}^{\ast}$ by simply rendering the CAD on the corresponding reference pose.
Considering the category-level setting, methods such as E-PnP [47] cannot be used; therefore, a new approach is needed to refine orientation and position transformations.
@@
We optimize for both rotation and translation $\mathbf{R}$ and $\mathbf{t}$. Given the rotated 3D reference points at each time step $k$ as $\gamma=\left(\mathbf{R}\mathbf{X}_{r}^{k}+\mathbf{t}\right)$, we minimize the error between the 2D reprojected reference points $\mathbf{K}_{q}\gamma$ (using query intrinsic matrix $\mathbf{K}_{q}$) and the matched 2D points $\mathbf{x}_{q}^{k}$ through:
$\left(\mathbf{R}_{q},\mathbf{t}_{q}\right)=\operatorname*{argmin}_{\mathbf{R},\mathbf{t}}\sum_{k=1}^{K}\left\lVert\mathbf{K}_{q}\gamma-\mathbf{x}_{q}^{k}\right\lVert_{2},$ (8)
where $\mathbf{R}_{q},\mathbf{t}_{q}$ are the resulting query rotation and translation respectively. This iterative refinement process leverages first-order stochastic gradients calculated with Adam [50] to gradually align points while optimizing for pose parameters.
@@ IV Experiments
The objective of the method is to show the possibility to perform category-level pose estimation from a single RGB query image and a set of few RGB references.
For this reason the translation vector computed in Eq.8 is obtained up-to an unknown scaling factor. To this end, the focus is on the rotation error (computed as the Geodesic distance on the $\text{SO}(3)$ manifold of rotation matrices). Such an evaluation is present in several RGB-based pose estimation works like POPE [34] and NOPE [52], which however are limited to instance-level scenarios. Similarly, the proposed evaluation is contained in ZSP in [8], which represents the baseline for quantitative comparison. Nevertheless, for real-world applications, when the query’s depth is available, the scaling factor can be obtained to opportunely rescale the translation vector coming from Eq.8. The assessment includes median rotation error (in degrees) and accuracy. The latter is computed as the percentage of estimates that fall below a certain rotation threshold. Here, 15 and 30 degrees thresholds have been chosen, following ZSP [8]. Results are provided both per category and average over all the categories. Additionally, qualitative results are presented on the Objectron dataset [53] to highlight the cross-dataset generalization ability of the proposed method to generalize to diverse RGB-only settings. All the experiments have been run on a single NVidia RTX3090 GPU. For each set of reference images, 50 novel views are generated as a trade-off between generation and 3D model reconstruction timing. The pose estimation optimizer runs 1000 iterations with an early stop on the loss computed as in Eq.8. This step takes 2-4 seconds to run, depending on the above stopping criteria. The most time-intensive step is the 3D CAD generation, which, however, is carried out only once per category, so in the long run, it does not influence the real-time factor of the application.
@@ IV-A1 Comparison against depth-free approach
Since the presented approach is RGB-only, it is worth starting a quantitative evaluation against methods employing no-depth images, which generally provide a great geometric prior and coherently improved performance. The concurrent work ZSP presented a depth-free version, based on the Essential Matrix calculation [8],[54]. The Essential Matrix embeds the relative pose information between two RGB views from a set of matches. It is typically used in stereo-vision, yet in this case has been effectively employed, by exploiting only the available sparse semantical correspondences.
@@ IV-A2 Comparison against RGB-D approach
The second mode, analyzed in Table II, encompasses the full method, where ZSP can show better performance only when enough views are available, by exploiting depth information for both query and reference. As it appears from Table II, with the diffusion-based novel-view synthesis, it is possible to substantially improve results over best view selection carried out in ZSP [8], thanks to having more images available for the matching phase. At the same time, this indicates the reduced sim-to-real gap between real and generated images, highlighting how diffusion models can be useful for this task. On the other hand, as shown in Table III, the Umeyama-RANSAC depth-based refinement method provided by ZSP increases the performance with respect to its correspondent best-view. This method requires an available and employable depth and boosts ZSP’s performance; however, it is eroded in few-view reference scenarios, as it is possible to notice a cross-over point with three views. Ultimately, Zero123-6D completes the overtake on ZSP when just a single reference view is available. When reducing from three views to one, Zero123-6D maintains a proper understanding of the 3D structure of the object. At the same time, the depth cues can no longer compensate for the small number of views, as in the previous cases. This can become relevant in fixed camera settings scenarios, where acquiring multiple views of the reference object is difficult.
@@ IV-B Qualitative results
This section highlights the qualitative results of the method in accordance with the aforementioned discussion about numerical performance. Fig.3 shows the feature map of the instances and the selected best-view. The best-view matching is possible thanks to the semantical information provided by DINO features. Applying Principal Component Analysis (PCA) on the feature map makes it possible to see how similar parts of the objects are depicted with similar colors. This is reflected in the quality of matches, which makes it possible to initialize the presented refiner. As further experiments, qualitative results on the Objectron [55] dataset are presented. These are particularly interesting because Objectron, unlike CO3D, does not have depth maps associated with RGB images. Hence, in this scenario, methods like ZSP [8], could not be effectively employed without a performance drop, while Zero123-6D’s method can be applied unaltered. Results presented in Fig.4 depict such cross-dataset generalization, representing a promising direction to employ objects generated through novel-view synthesis for pose estimation in more challenging setups (e.g. cluttered scenarios).